40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16837     Entry time: Mon May 9 18:43:03 2022     In reply to: 16832     Reply to this: 16838   16843
Author: Anchal 
Type: Update 
Category: BHD 
Subject: ITMX table layout corrected 

As I went to correct the ITMX Oplev mirrors, I found that both mirrors were placed in very different positions than the design position. Part of the reason I think was to preserve outside oplev path, and party because a counterweight was in ITMXOL1 position. I had to do following steps to correct this:

  • I noted down level meter readings of the table before making any changes.
  • I removed the counter weight from near the center of the table.
  • I placed the Oplev mirrors in the nominal positions.
  • I placed the counter weight near previous position.
  • I moved a edge hanging counter weight to get back the level meter to its previous state coarsely.
  • Then I used dataviewer to find the previous OSEM PD monitor values and changed ITMX PIT and YAW to come closer to those PD values. And voila, I regained the flashing on Xarm. I nudged the ITMX pit and yaw bit more to maximize it.
  • I then went back to aligning the Oplevs properly.
  • Then I adjusted the POP mirrors to get the beam back through center of window. This was very tricky and took a lot of time.
  • Now the beam is going through near center and the oplev beams are far away enough from POP_SM5.
  • On the outside table, I noted the POP beam and the oplev beam. I corrected the pit of the returning beam to get the oplev beam at nominal height on outside table.

ITMX Sat Amp is flaky

[Anchal, Paco]

During the above work, i must have kicked the cable between the vacuum flange and the satellite amplifier box for ITMX. This disconnected all the OSEMs and Coils. We tried several things to debug this and finally found that nudging the connections on Sat Amp box brought the OSEMs and coils back online. Note that the connector was not partially out or in a state that obviously showed disconnection of the pins. I'm glad we are putting in new electronics soon for the vertex optics as well.


Next steps:

  • I showed Tega the returning oplev beam and the POP beam coming out of the ITMX chamber.
  • The Oplev beam paths need to be adjusted.
    • The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
    • First setup two irises to save ingoing path.
    • Then make space for the returning beam by changing the steering mirror positions.
    • Then recover the ingoing path to the center of irises.
    • Steer the returning beam to the QPD.
    • Then maximize the flashing on XARM and center the oplev to save this position.
  • POP beam needs to be directed to previous setup on far side of table.
    • The POP beam is coming out at the rising angle.
    • This is good for us if we do bit unconventional stuff and transfer the beam to other side of table at an elevated height. Given how close all the beams are coming out of the viewport, I think this is the best solution in terms of saving time.
    • Get the beam down to the old setup which was camera and photodiodes all aligned.

 

 

ELOG V3.1.3-